OpenStudy (anonymous):

Find the coordinates of the midpoint of AB. A(12, -7) and B(-4, 7). A. (-8, 7) B. (4,0) C. (8, -7) D. (8, 0)

OpenStudy (jdoe0001):

\(\large {\begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&(12\quad ,&-7)\quad B&(-4\quad ,&7) \end{array}\\ \quad \\ \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)}\)

OpenStudy (anonymous):

So im suppose to do 12+-4 & -7+7. for those i got 8 and 0. is that correct?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (ranga):

Don't forget to divide by 2.

OpenStudy (ranga):

(12 - 4) / 2 = ? (-7 + 7) / 2 = ?

OpenStudy (anonymous):

ok thanks for helping! means alot!(:

OpenStudy (ranga):

you are welcome.